OpenStudy (janu16):

Erika is working on solving the exponential equation 50x = 17; however, she is not quite sure where to start. Using complete sentences, describe to Erika how to solve this equation.

OpenStudy (janu16):

@phi

OpenStudy (phi):

you should use ^ to show exponents. like so: 50^x = 17 or the equation editor \[ 50^x = 17 \]

OpenStudy (janu16):

its 50^x by the way

OpenStudy (janu16):

yea sorry

OpenStudy (phi):

exponents and logs go together , so you have to remember that (they "undo" the other) here is the rule you want to remember: \[ \log\left( a^b\right)= b \log(a) \]

OpenStudy (janu16):

ok

OpenStudy (janu16):

so do i put that it the form you showed?

OpenStudy (phi):

you have an equation. if you do the "same thing" to both sides it will stay equal in this case, "take the log" of both sides for example , if you had 2^x = 8 take the log means write \[ \log(2^x) = \log(8) \]

OpenStudy (janu16):

log(50^x)=log(17)?

OpenStudy (phi):

exactly. now on the left side, we use the "rule" (or property of logarithms) \[ \log\left( a^b\right)= b \log(a) \] to rewrite the left side can you do that ?

OpenStudy (janu16):

\[x \log 50 = \log 17\]

OpenStudy (janu16):

?

OpenStudy (phi):

yes. looks good. even though it looks complicated, log(50) is a number we can divide both sides by the same number i.e. by log(50) \[ x \frac{\log(50)}{\log(50)} = \frac{\log(17)}{\log(50)}\] on the left side log50 divided by itself is 1 we get \[ x = \frac{\log(17)}{\log(50)}\]

OpenStudy (janu16):

so thats the final answer?

OpenStudy (phi):

you could use a calculator, and find a decimal value. but yes, that is the answer. x is about 0.72423228 type into google: 50^0.72423228 = and you should get something close to 17